OpenStudy (anonymous):

PLEASE HELP ME Solve by the linear combination method (with or without multiplication). 2x + 3y = –17 5x + 2y = –4 (3, –1) (–2, 2) (–1, 3) (2, –7)

OpenStudy (anonymous):

okay need help

OpenStudy (anonymous):

(2, -7)

OpenStudy (anonymous):

a) |x + y = 5 |2x - y = 7; b) |2x + y = 5 |x - y = 2 c) |3x + y = 6 |4x - 3y = -5 d) |1/(x - 1) = y - 3 |x - y = -2 e) |(9x + 4y)/3 - (5x - 11)/2 = 13 - y |13x - 7y = -8 Answer: c and e has solution (1; 3) Example

OpenStudy (anonymous):

a or b

OpenStudy (quin100):

do you need a drawing of it?

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

how

OpenStudy (anonymous):

you say a or b and they say c or d

OpenStudy (anonymous):

ow

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

why i think a and b 4 is -17 is not in one of the answers a , b d

OpenStudy (godlovesme):

its (2,-7) if u plug 2 for x and -7 for y it will make ur eqns true does that help @welligya...

OpenStudy (quin100):

A

OpenStudy (anonymous):

thanks

OpenStudy (godlovesme):

thanks for the medal :) and anytime ^.^

OpenStudy (anonymous):

A

OpenStudy (quin100):

you welcome

OpenStudy (godlovesme):

@Mesopretty and @quin100 its D. (2,-7)